A body of mass 400 g slides on a rough horizontal surface. If the frictional force is 3.0 N, find 
  1. the angle made by the contact force on the body with the vertical .
  2. the magnitude of the contact force.

Grade:12th pass

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
The Normal Reaction would be given by:
N=mg=0.400\times 10=4 N
The frictional force is given by:
f=3 N
Hence the angle made by the contact force would be given by:
\theta =tan^{-1}\left (\frac{3}{4} \right )=37^{0}
and the magnitude of the contact force is given by:
F=\sqrt{3^{2}+4^{2}}=\sqrt{9+16}=5 N
Regards
Sumit